What are the values of x and y for the following reaction if E∘=0.91V and ΔG∘=−527kJ?

3Ax++2B→3A+2By+
Answer #1

\Delta G^{o}  = - nFEo

given, \Delta G^{o} = -527 KJ = -527000 J

Eo = 0.91 V

F = 96500 C

putting the given values

- 527000 = - n × 96500× 0.91

Or, n = (527000/96500×0.91)

Or, n = 6.

So, total electron transferred = 6.

Total postive charge on left side = total postive charge on right side = 6

Therefore,

Or, 3×x = 6

Or, x = (6/3) = 2

and,

2y = 6

Or, y = (6/2) = 3.

Final equation is

3A2+ + 2B \to 3A + 2B3+